OpenStudy (anonymous):

X^2 - 2x - 13 = 0 using the India method for solving quadratic equations

OpenStudy (anonymous):

x^2 -2x +1 =14 (x-1)^2=14

OpenStudy (anonymous):

x-1 = = + or - sqrt(14)

OpenStudy (anonymous):

x= 1 +sqrt(14) or 1 - sqrt(14)

OpenStudy (anonymous):

x^2 - 2x - 13 = 0 x^2 - 2x =13 4x^2 - 8x = 52 4x^2- 8x + 4 = 56 (2x - 2)^2= 56 2x - 2 = √(56) or 2x - 2 = -√(56) x = (√(56)+2)/2 or x = (-(√56)+2)/2

OpenStudy (anonymous):

what is the india method

OpenStudy (anonymous):

sounds made up

OpenStudy (anonymous):

no....actually some Indian solved d quadratic equation by making complete square in ancient time....do google...u'll come to know.

OpenStudy (anonymous):

even I did wrong...there is certain procedure for completing the square